# The art of modeling nuclear reactions with weakly bound nuclei: status and perspectives

### v1

@@ 7.6 Coulomb dissociation revisited: quantum-mechanical and nuclear effects
To overcome (i), two alternative strategies have been commonly adopted. The first consists in measuring the breakup data at sufficiently small c.m. scattering angles, under the assumption that, for angles well below the grazing one, nuclear effects are negligible (the so-called “safe-Coulomb” region). This method has been adopted, for example, in exclusive breakup experiments performed at RIKEN, at typical energies around 100 MeV per nucleon (e.g. Nak99 ; Nak06 ; Fuk04 ). A second, alternative, strategy is to collect the breakup data in a wider angular range. This has the advantage of increasing the statistics, but the disadvantage of including nuclear contributions. The latter must be estimated and subtracted from the breakup data for a meaningful extraction of the electric contribution.
@@
Each of these strategies has pros and cons. The “safe-Coulomb method” has the disadvantage of considering only a fraction of the breakup data. Moreover, nuclear breakup can extend down to very small angles. For example, in the analysis of the 11Li+208Pb Coulomb dissociation experiment performed at RIKEN at an incident energy of 70 MeV per nucleon Nak06 , the authors considered data below $1.46^{\circ}$.
@@
The “safe-Coulomb” assumption is tested in Fig. 55 for the 11Be +208Pb reaction, which shows the n-10Be relative energy distribution for two upper cuts of the scattering angle, $\theta_{\mathrm{max}}=6^{\circ}$ (top) and $1.3^{\circ}$ (bottom). The lines are XCDCC calculations based on a particle-plus-rotor model of 11Be, including both Coulomb and nuclear couplings or only nuclear couplings. It is evident from the figure that nuclear effects are still very relevant for $\theta_{\mathrm{max}}=6^{\circ}$ but are almost suppressed (but not negligible) for the lower cut. This gives an idea of the accuracy and validity of the EPM analysis performed in Fuk04 to extract the $B(E1)$ distribution for 11Be.
@@
The alternative strategy, in which angle-integrated cross sections are considered for the analysis, has the advantage of profiting from the entire breakup data, but deals with the uncertainties associated with the estimation of the nuclear contribution. For that, measurements with a high-$Z$ target nucleus are combined with those for a light nucleus, like 12C, for which the nuclear contribution is expected to be dominant. Thus, the nuclear contribution in the reaction with the heavy target is accounted for by subtracting a properly scaled cross section measured with the carbon target. The scaling factor can be estimated from the respective radii of the heavy and light targets. Alternatively, one may resort to eikonal formulas for stripping and diffraction to estimate the cross section ratio between heavy and light targets, as done in Pal03 .
Besides the additional source of uncertainty added by the need of estimating the nuclear contribution, another caveat of this method is the neglect of possible interference effects between the nuclear and Coulomb contributions. The importance of these effects can be assessed by means of quantum-mechanical calculations, such as those given by the CDCC method.
As an example, in Fig. 56 the effect of Coulomb/nuclear interference is illustrated for the 11Be+208Pb reaction at 520 MeV/u. The plot shows EPM and XCDCC calculation. In the latter, different choices for the nuclear and Coulomb couplings are shown. It is seen that (i) the nuclear breakup is rather significant (consistently with the estimate of Pal03 ) and, more importantly, that (ii) the incoherent sum of the nuclear and Coulomb contributions overestimates the full XCDCC calculations in which Coulomb and nuclear couplings are included simultaneously and hence interference effects appear. From these calculations, it becomes apparent that the procedure of summing incoherently the nuclear and Coulomb contributions introduces some error in the extraction of the electric transition probabilities from Coulomb dissociation experiments.
@@
In addition of nuclear contribution and its interference with the Coulomb contribution, higher-order couplings effects may also affect the extraction of electric transition probabilities from cross section data. In principle, one might rely on more sophisticated calculations, such as those provided by the CDCC and XCDCC methods, in which nuclear and Coulomb couplings are included to all orders and their interference is properly accounted for. The price to pay is that these calculations are more computationally demanding, and more importantly, they do not provide a simple connection between the reaction observables (cross sections) and the underlying $B(E\lambda)$ distributions. A recent work Mor20 proposed a method to extract $B(E1)$ distribution by comparing exclusive breakup data with CDCC calculations based on some reasonable trial structure model of the weakly-bound nucleus, which is subsequently corrected by comparing the CDCC and experimental cross sections.
@@
An application of this method is shown in Fig. 57, adapted from Mor20 . The solid black line is the $B(E1)$ distribution provided by the starting $B(E1)$ model used in the XCDCC calculation. This $B(E1)$ is corrected for each excitation energy by comparing the calculated and measured cross sections, as explained in Mor20 . The procedure is applied separately to the data of Ref. Fuk04 , for the two angular cuts discussed above, as well as to the data of Pal03 . The corrected $B(E1)$ distributions are displayed in this figure with the symbols. It is seen that the three of them agree reasonably well within the estimated errors. This is in contrast to the $B(E1)$ extracted in the original works and shown in Fig. 48. It is important to note that the latter are affected by the experimental resolution, while those shown in Fig. 57 do not include the effect of the experimental energy resolution. Differences between these experimental resolutions in the two experiments are partially responsible for the differences observed in the $B(E1)$ distributions reported in the original works.
@@
Figure 57: Unfolded $B(E1)$ distributions extracted from the experimental breakup cross sections measured in Pal03 (squares) and Fuk04 (circles) using the procedure based on the XCDCC method proposed in Mor20 . For comparison, the starting $B(E1)$ distribution is shown by the black line and NCSMC ab-initio calculation of Ref. Cal16 (red solid line) is also shown. Figure adapted from Mor20 .
@@ 8 Study of nucleon-nucleon correlations from nucleon removal reactions
One particular point of interest for nucleon-Borromean systems of the form N+N+C is the study of the correlation between their two valence nucleons, since the nucleon-nucleon interaction is essential to hold the Borromean system bound. As many Borromean systems also present nucleon halos, the study of the correlation between their valence nucleons permits the exploration of the nucleon-nucleon interaction in low-density, isospin-asymmetric environments. A quantity related to the nucleon-nucleon correlation is the average opening angle between nucleons with respect to the core $\left\langle\theta_{NN}\right\rangle$. If both nucleons follow independent uncorrelated orbits, this angle will on average be $90^{\circ}$ Esb92 ; Ber07 . A smaller angle corresponds to a small average distance between nucleons compared to the distance of their center of mass to the core, in which is known as a “dineutron” (or diproton) configuration mig73 , which is associated with an attractive correlation between nucleons. Meanwhile, an angle larger than 90 indicates a larger internucleon distance compared to their distance to the core, which corresponds to both nucleons being on opposite sides of the core, in the so-called “cigar” configuration, which corresponds to a repulsive nucleon-nucleon correlation.
@@
where $\left\langle r_{n}^{2}\right\rangle$ is the average square distance between neutron and core, $Z_{c}$ is the charge of the core and $A$ the mass of the nucleus Bert91 , so an enhanced $B(E1)$ indicates a smaller angle and dineutron correlations while a reduced $B(E1)$ indicates the existence of a cigar-like configuration. $B(E1)$ strengths have been measured for various Borromean nuclei, from which values for $\left\langle\theta_{NN}\right\rangle$ can be extracted. The values extracted in Ber07 are presented in Table 3, where the angles can be seen to suffer from significant uncertainties, due both to experimental uncertainties and the model dependence due to $\left\langle r_{n}^{2}\right\rangle$.
@@
Another method of extracting the opening angle is the use of nucleon knockout reactions, where one of the valence nucleons is removed via a sudden reaction, such as nucleon-knockout with 9Be or 12C targets or a $(p,pN)$ reaction with proton targets, assuming the validity of the spectator approximation for the remaining $(C+N)$ system. Within this approximation, in the projectile rest frame the Borromean nucleus start at rest, so that, when the nucleon is removed, the final momentum of the $(C+N)$ system will be equal and opposite to that of the removed nucleon within the Borromean system ($\boldsymbol{k}_{y}$). Since the $(C+N)$ system is unbound, the relative momentum between $C$ and the remaining nucleon can also be measured $(\boldsymbol{k}_{x})$, and the angle between them $\theta_{k}$ determined. A scheme of the quantities of interest is shown in Fig. 58.
@@
It should be noted that, as $\theta_{k}$ is an angle between the momenta associated to the nucleons instead of their positions, it can be viewed as a Fourier transform of the opening angle $\theta_{NN}$. As such, a value of $\theta_{k}$ larger than $90^{\circ}$ is associated to a value of $\theta_{NN}$ lower than $90^{\circ}$, and thus to a dineutron configuration, while the cigar configuration corresponds to $\theta_{k}<90^{\circ}$ and $\theta_{NN}>90^{\circ}$. Various measurements of $\theta_{k}$ have been performed for Borromean nuclei Chu97 ; Sim99 ; Mar01 ; Sim07 , finding for its average value $\left\langle\theta_{k}\right\rangle$ the following results (in degrees): $90$ (6He), $90$ (8He), $103.6^{+0.7}_{-0.9}$ (11Li), $96.2^{+1.8}_{-1.6}$ (14Be) Sim07 . It can be seen that, indeed, values for $\left\langle\theta_{k}\right\rangle>90^{\circ}$ correspond to $\left\langle\theta_{NN}\right\rangle<90^{\circ}$, as shown in Table 3. It should also be noted that an asymmetric distribution in either $\theta_{k}$ or $\theta_{NN}$ (resulting in $\left\langle\theta\right\rangle\neq 90^{\circ}$) requires the interference of waves with opposite parities Cat84 , which neatly explains the symmetry for 6,8He, since the valence neutrons find themselves in the negative-parity $p_{3/2}$ and $p_{1/2}$ waves, so their angular distribution must be symmetric. This also sets the average opening angle as a measure of the admixture of opposite parity components Kub20 .
@@
where $\boldsymbol{x}$ and $\boldsymbol{y}$ are defined as in Sec. 4.2, $\phi_{gs}$ describes the bound state of the Borromean nucleus, $\phi_{Cn}$ the final-state wave-function between the core and the remaining neutron, and $S(y)$ is a $S-$matrix that includes the absorptive potential between target proton and core $C$. As such, the cross section can be interpreted as the square of the Fourier transform of the bound state, distorted by the absorption of the target proton and the final-state interaction between neutron and core. In Kub20 , the maximum $\left\langle\theta_{k}\right\rangle$ for 11Li was found for small values of $k_{y}\sim 0.3$ fm-1, which was interpreted as the dineutron being located at the nuclear surface where the nuclear density is low, as predicted by previous models Mats06 ; Hag07 . As the low nuclear density is a feature of the surface of all neutron-Borromean nuclei, it is expected that the dineutron also appears in the surface of other such systems. Indeed, in Cor23 , the maximum of $\left\langle\theta_{k}\right\rangle$ was found for roughly the same value of $k_{y}$ also for 14Be and 17B, as shown in Fig. 59, making a tantalizing case for the universality of dineutron correlations in the surface of neutron-Borromean nuclei. as predicted in Mats06 , although experiments on more nuclei, such as 19B or 22C, are required in order to confirm this feature. In the figure the green bands correspond to systematic errors in experimental data, while statistical errors are indicated by the errorbars. The theoretical calculations are able to match experimental results for 11Li and 17B but overestimate the maximum average angle for 14Be. This has been associated to the contribution of excited states of the 12Be beyond the $2^{+}$ state included in the theoretical model Cor23 .

### v2

@@ 8.1 Semiclassical analysis of Coulomb dissociation data
To overcome this difficulty, two alternative strategies have been adopted. The first consists in measuring the breakup data at sufficiently small c.m. scattering angles, under the assumption that, for angles well below the grazing one ($\theta_{\mathrm{gr}}$), nuclear effects are negligible (the so-called “safe-Coulomb” region). This method has been adopted, for example, in several exclusive breakup experiments performed at RIKEN, at typical energies of around 100 MeV per nucleon (e.g. Nak99 ; Nak06 ; Fuk04 ). A second, alternative, strategy is to collect the breakup data in a wider angular range. This has the advantage of increasing the statistics, but the disadvantage of including nuclear contributions. The latter must be estimated and subtracted from the breakup data for a meaningful extraction of the electric contribution.
@@
The “safe-Coulomb method” has the disadvantage of considering only a fraction of the breakup data. In fact, nuclear breakup can extend down to very small angles. For example, in the analysis of the 11Li+208Pb Coulomb dissociation experiment performed at RIKEN at an incident energy of 70 MeV per nucleon Nak06 , the authors considered data below $1.46^{\circ}$. The alternative method of considering a wider angular range has the obvious challenge of requiring a realistic method to estimate the nuclear contribution. In the remainder of this section we discuss further these two methods.
For the practical application of the “safe-Coulomb” method, one integrates the double differential cross section given by Eq. (134) up to a maximum angle ($\theta_{\mathrm{max}}$) such that for $\theta<\theta_{\mathrm{max}}$ breakup can be assumed to be purely Coulomb and, therefore, (134) is valid ($\theta_{\mathrm{max}}<\theta_{\text{gr}}$, with $\theta_{\text{gr}}$ the grazing angle). This gives:
[…]
Restricting the analysis to the lower cut ($\theta_{\mathrm{c.m.}}<1.3^{\circ}$), and assuming that for these angles the dissociation probability is dominated by the Coulomb dipole couplings, one can use the approximation in Eq. (154) and extract the $B(E1$) distribution. The result of such an analysis is shown in Fig. 58, which displays the $B(E1)$ extracted in Fuk04 (circles).
The alternative strategy, in which angle-integrated cross sections are considered for the analysis, has the advantage of profiting from the entire breakup data, but deals with the uncertainties associated with the estimation of the nuclear contribution. For that, measurements with a high-$Z$ target nucleus are combined with those for a light nucleus, like 12C, for which the nuclear contribution is expected to be dominant. Thus, the nuclear contribution in the reaction with the heavy target is accounted for by subtracting a properly scaled cross section measured with the light target. The scaling factor can be estimated from the respective radii of the heavy and light targets. Alternatively, one may resort to eikonal formulas for stripping and diffraction to estimate the cross section ratio between heavy and light targets, as done in Pal03 .
In addition to the additional source of uncertainty caused by the need to estimate the nuclear contribution, another caveat of this method is the neglect of possible interference effects between nuclear and Coulomb contributions. The importance of these effects can be assessed by means of quantum-mechanical calculations, such as those given by the CDCC method.
If the approximations implied by the EPM are well fulfilled, the extracted $B(E\lambda$) distributions should be independent of the incident energy and the target as well as of the adopted analysis method. A consistency test can be found in the $B(E1)$ distributions extracted for the 11Be nucleus obtained from the analysis of two 11Be+208Pb experiments performed at GSI by Palit et al. Pal03 at 70 MeV/u and 520 MeV/u, shown in Fig. 58. (The origin of this discrepancy is further discussed below). The two distributions are compatible but with clear differences at the peak.
8.2 Quantum-mechanical effects: CDCC analysis of Coulomb dissociation data
As an alternative to the semiclassical methods discussed in the previous section, one can use quantum-mechanical methods, such as CDCC, discussed in Sec. 3. In CDCC, nuclear and Coulomb breakup are included on an equal footing to all orders, and their interference is properly taken into account.
@@ 8.2 Quantum-mechanical effects: CDCC analysis of Coulomb dissociation data
In addition, the CDCC method can be used to assess the validity of the approximations involved in the semiclassical approaches. For example, the “safe-Coulomb” assumption is tested in Fig. 59 for the 11Be +208Pb reaction, which shows the n-10Be relative energy distribution for two upper cuts of the scattering angle, $\theta_{\mathrm{max}}=6^{\circ}$ (top) and $1.3^{\circ}$ (bottom). The lines are XCDCC calculations based on a particle-plus-rotor model of 11Be, including both Coulomb and nuclear couplings or only nuclear couplings. It is evident from the figure that nuclear effects are still very relevant for $\theta_{\mathrm{max}}=6^{\circ}$ but are almost suppressed (but not negligible) for the lower cut. This gives an idea of the accuracy and validity of the EPM analysis performed in Fuk04 to extract the $B(E1)$ distribution for 11Be.
@@
The CDCC method can be also used to test the importance of Coulomb/nuclear interference effects. This is illustrated in Fig. 60 for the 11Be+208Pb reaction at 520 MeV/u. The plot shows EPM and XCDCC calculation. In the latter, different choices for the nuclear and Coulomb couplings are shown. It is seen that (i) nuclear breakup is rather significant (consistently with the estimate of Pal03 ) and, more importantly, (ii) the incoherent sum of the nuclear and Coulomb contributions overestimates the full XCDCC calculations with interference effects. From these calculations, it becomes apparent that the procedure of summing incoherently the nuclear and Coulomb contributions introduces some error in the extraction of the electric transition probabilities from Coulomb dissociation experiments. The error will become larger at lower incident energies.
@@
Ideally, one may also analyze directly the experimental cross sections with the CDCC method to extract the electric transition probabilities and other structure properties. The practical applicability of this idea is hindered by the fact that there is no simple connection between the reaction observables (cross sections) and the underlying $B(E\lambda)$ distributions. In particular, breakup cross sections are no longer proportional to the $B(E\lambda)$.
One possibility is to compare the measured cross sections with a series of CDCC calculations assuming different structure models, with different electric strength distributions. This is exemplified in Fig. 61 where the data correspond to the relative energy distribution of 15C on 208Pb at 68 MeV/u measured at RIKEN Nak03 and the lines are CDCC calculations for different potential models of the 15C nucleus. From the comparison one may select the best structure model among those considered.
Figure 61: Comparison of the Coulomb dissociation data of 15C on 208Pb at 68 MeV/u Nak03 with CDCC calculations using different 14C+n models for 15C. Figure from Sum08 .
Recently, an alternative to this procedure to extract $B(E\lambda)$ distributions from the comparison of CDCC calculations with breakup cross section data has been proposed Mor20 . The strategy can be summarized as follows:
  • Start with some trial structure model of the projectile. This will predict some distribution $B^{0}(E1;\varepsilon)$ depending on the continuum energy.
  • Introduce some small changes (perturbation) in the model, by multiplying all Coulomb dipole matrix elements by arbitrary factors $(1+2\delta(\varepsilon_{i}))$, where $\varepsilon_{i}$ are the measured energies. This will modify the $B(E1)$ at each measured $\varepsilon_{i}$
    $B^{\text{mod}}(E1,\varepsilon_{i})\simeq B^{0}(E1,\varepsilon_{i})\left(1+2\delta(\varepsilon_{i})\right),$ (155)
  • For small perturbations, the calculated breakup cross sections are modified simply as Mor20
    $\sigma^{\text{mod}}_{i}\simeq\sigma^{0}_{i}+\delta(\varepsilon_{i})\,{\sigma^{\prime}}_{i}.$ (156)
    where $\sigma^{0}_{i}$ and ${\sigma^{\prime}}_{i}$ are constants which can be determined by performing several calculations for different $\delta$ values.
  • By comparing the model calculations with the measured breakup data, one may infer the optimal value of $\delta(\varepsilon_{i})$ for each excitation energy and, from them, the values of $B(E1;\varepsilon^{i})$ that best describe the data within the CDCC framework.
@@
An application of this method is shown in Fig. 62, adapted from Mor20 . The solid black line is the $B(E1)$ distribution provided by the starting (trial) $B(E1)$ model used in a XCDCC calculation. Following the outlined procedure, this $B(E1)$ is corrected for each excitation energy by comparing the calculated and measured cross sections. The procedure is applied separately to the data of Ref. Fuk04 , for the two angular cuts discussed above, as well as to the data of Pal03 . The corrected $B(E1)$ distributions are displayed in this figure with the symbols. It is seen that the three of them agree reasonably well within the estimated errors. This is in contrast to the $B(E1)$ extracted in the original works and shown in Fig. 58. It is important to note that the latter are affected by the experimental resolution, while those shown in Fig. 62 do not include the effect of the experimental energy resolution. Differences between these experimental resolutions in the two experiments are partially responsible for the differences observed in the $B(E1)$ distributions reported in the original works.
@@
Figure 62: Unfolded $B(E1)$ distributions extracted from the experimental breakup cross sections measured in Pal03 (squares) and Fuk04 (circles) using the procedure based on the XCDCC method proposed in Mor20 . For comparison, the starting $B(E1)$ distribution is shown by the black line and a NCSMC ab-initio calculation of Ref. Cal16 (red solid line) is also shown. Figure adapted from Mor20 .
@@ 9 Study of nucleon-nucleon correlations from nucleon removal reactions
One particular point of interest for nucleon-Borromean systems of the form N+N+C is the study of the correlation between their two valence nucleons, since the nucleon-nucleon interaction is essential to hold the Borromean system bound. As many Borromean systems also present nucleon halos, the study of the correlation between their valence nucleons permits the exploration of the nucleon-nucleon interaction in low-density, isospin-asymmetric environments. A quantity related to the nucleon-nucleon correlation is the average opening angle between nucleons with respect to the core $\left\langle\theta_{NN}\right\rangle$. If both nucleons follow independent uncorrelated orbits, this angle will on average be $90^{\circ}$ Esb92 ; Ber07 . A smaller angle corresponds to a small average distance between nucleons compared to the distance of their center of mass to the core, in which is known as a “dineutron” (or diproton) configuration mig73 , which is associated with an attractive correlation between nucleons. Meanwhile, an angle larger than 90 indicates a larger internucleon distance compared to their distance to the core, which corresponds to both nucleons being on opposite sides of the core, in the so-called “cigar” configuration, which corresponds to a repulsive nucleon-nucleon correlation. Note that three-body calculations show that both “dineutron” and “cigar” components appear in the wave function Zhu93 , and it is their relative strength which leads to a more dominant configuration.
@@
where $\left\langle r_{n}^{2}\right\rangle$ is the mean square distance between neutron and core, $Z_{c}$ is the charge of the core and $A$ the mass of the nucleus Bert91 , so an enhanced $B(E1)$ indicates a smaller angle and dineutron correlations while a reduced $B(E1)$ indicates the existence of a cigar-like configuration. $B(E1)$ strengths have been measured for various Borromean nuclei, from which values for $\left\langle\theta_{NN}\right\rangle$ can be extracted. The values extracted in Ber07 are presented in Table 3, where the angles can be seen to suffer from significant uncertainties, due both to experimental uncertainties and the model dependence due to $\left\langle r_{n}^{2}\right\rangle$.
@@
Another method of extracting the opening angle is the use of nucleon knockout reactions, where one of the valence nucleons is removed via a sudden reaction, such as nucleon-knockout with 9Be or 12C targets or a $(p,pN)$ reaction with proton targets, assuming the validity of the spectator approximation for the remaining $(C+N)$ system. Within this approximation, in the projectile rest frame the Borromean nucleus starts at rest, so that when the nucleon is removed, the final momentum of the $(C+N)$ system will be equal and opposite to that of the removed nucleon within the Borromean system ($\boldsymbol{k}_{y}$). Since the $(C+N)$ system is unbound, the relative momentum between $C$ and the remaining nucleon can also be measured $(\boldsymbol{k}_{x})$, and the angle between them $\theta_{k}$ determined. A scheme of the quantities of interest is shown in Fig. 63.
@@
It should be noted that, as $\theta_{k}$ is an angle between the momenta associated with the nucleons instead of their positions, it can be viewed as a Fourier transform of the opening angle $\theta_{NN}$. As such, a value of $\theta_{k}$ larger than $90^{\circ}$ is associated with a value of $\theta_{NN}$ lower than $90^{\circ}$, and thus to a dineutron configuration, while the cigar configuration corresponds to $\theta_{k}<90^{\circ}$ and $\theta_{NN}>90^{\circ}$. Various measurements of $\theta_{k}$ have been performed for Borromean nuclei Chu97 ; Sim99 ; Mar01 ; Sim07 , finding for its average value $\left\langle\theta_{k}\right\rangle$ the following results (in degrees): $90$ (6He), $90$ (8He), $103.6^{+0.7}_{-0.9}$ (11Li), $96.2^{+1.8}_{-1.6}$ (14Be) Sim07 . It can be seen that, indeed, values for $\left\langle\theta_{k}\right\rangle>90^{\circ}$ correspond to $\left\langle\theta_{NN}\right\rangle<90^{\circ}$, as shown in Table 3. It should also be noted that an asymmetric distribution in either $\theta_{k}$ or $\theta_{NN}$ (resulting in $\left\langle\theta\right\rangle\neq 90^{\circ}$) requires the interference of waves with opposite parities Cat84 , which neatly explains the symmetry for 6,8He, since the valence neutrons find themselves in the negative-parity $p_{3/2}$ and $p_{1/2}$ waves, so their angular distribution must be symmetric. This also sets the average opening angle as a measure of the admixture of opposite parity components Kub20 .
@@
where $\boldsymbol{x}$ and $\boldsymbol{y}$ are defined as in Sec. 4.2, $\phi_{gs}$ describes the bound state of the Borromean nucleus, $\phi_{Cn}$ the final-state wave-function between the core and the remaining neutron, and $S(y)$ is a $S-$matrix that includes the absorptive potential between target proton and core $C$. As such, the cross section can be interpreted as the square of the Fourier transform of the bound state, distorted by the absorption of the target proton and the final-state interaction between neutron and core. In Kub20 , the maximum $\left\langle\theta_{k}\right\rangle$ for 11Li was found for small values of $k_{y}\sim 0.3$ fm-1, which was interpreted as the dineutron being located at the nuclear surface where the nuclear density is low, as predicted by previous models Mats06 ; Hag07 . As the low nuclear density is a feature of the surface of all neutron-Borromean nuclei, it is expected that the dineutron also appears in the surface of other such systems. Indeed, in Cor23 , the maximum of $\left\langle\theta_{k}\right\rangle$ was found for roughly the same value of $k_{y}$ also for 14Be and 17B, as shown in Fig. 64, making a tantalizing case for the universality of dineutron correlations in the surface of neutron-Borromean nuclei, as predicted in Mats06 , although experiments on more nuclei, such as 19B or 22C, are required in order to confirm this feature. In the figure the green bands correspond to systematic errors in experimental data, while statistical errors are indicated by the error bars. The theoretical calculations are able to match experimental results for 11Li and 17B but overestimate the maximum average angle for 14Be. This has been associated with the contribution of excited states of the 12Be beyond the $2^{+}$ state included in the theoretical model Cor23 .